# The Buffer Mechanism for Multi-Step Information Reasoning in Language Models

### v2

@@ Abstract
Large language models have consistently struggled with complex reasoning tasks, such as mathematical problem-solving. Investigating the internal reasoning mechanisms of these models can help us design better model architectures and training strategies, ultimately enhancing their reasoning capability. In this study, we constructed a symbolic dataset to investigate the mechanisms by which Transformer models employ vertical thinking strategy based on their inherent structure and horizontal thinking strategy based on Chain of Thought to achieve multi-step reasoning. We introduced the concept of buffer mechanism: the model stores various information in distinct buffers and selectively extracts them through the query-key matrix. We proposed a random matrix-based algorithm to enhance the model’s reasoning ability, resulting in a 75% reduction in the training time required for the GPT-2 model to achieve generalization capability on the PrOntoQA dataset. These findings provide new insights into understanding the mechanisms of large language models.
@@ 1 Introduction
In recent years, LLMs have emerged and demonstrated remarkable capability across various tasks (Vaswani et al., 2017; Liu et al., 2018; Devlin et al., 2018; Radford et al., 2019; Touvron et al., 2023; OpenAI, 2023). These models have shown impressive in-context learning abilities (Brown et al., 2020; Dong et al., 2022; Garg et al., 2022) and have been applied to logical reasoning problems, such as matching top human contestants at the International Mathematical Olympiad (IMO) level (Trinh et al., 2024) and solving math problems (Davies et al., 2021).
However, even the most advanced models still struggle with complex reasoning tasks, which indicates the ability of LLMs to handle multi-step logical reasoning remains constrained. To truly enhance the reasoning capability of LLMs, it is crucial to investigate their intrinsic mechanisms.
Multi-step reasoning tasks encompass a broad concept, typically referring to the ability of a model to synthesize numerous complex conditions to answer questions. Here, we consider a representative class of syntactic structures within multi-step reasoning tasks: a sentence that includes both the question and sufficient known information to answer it. For example, “Given: [A]$\to$[B]...[B]$\to$[C]..., Question: 2-step reasoning starting from [A]", where “..." represents other textual content unrelated to logical reasoning. The answer to this question is “[C]". When a sentence contains only one logical reasoning step, it is often handled by the so-called induction head in Transformer (Brown et al., 2020; Olsson et al., 2022). However, multi-step reasoning is not merely a linear accumulation of multiple induction heads but involves more complex mechanisms.
Figure 1: Failure cases of multi-step reasoning in LLM. The correct answer is [r] (Since [w] $\to$ [p] $\to$ [e] $\to$ [i] $\to$ [r]). We tested each model 9 times.
@@
Large models employ two primary strategies for logical reasoning. The first, known as the Vertical Thinking Strategy (VTS), outputs reasoning results in a single forward based on their inherent structure. This approach is efficient but demands a high level of intelligence from the model itself. As shown in Fig. 1, current large models exhibit significant limitations in their vertical thinking capability. Another relatively less efficient approach is the Horizontal Thinking Strategy (HTS), such as Chain of Thought (CoT) (Wei et al., 2022; Kojima et al., 2022) and Tree of Thought (ToT) (Yao et al., 2024), Diagram of Thought (DoT) (Zhang et al., 2024a). This strategy substantially enhances the model’s reasoning performance. All models can produce the correct answers for tasks depicted in Fig. 1 with the help of CoT. Considering the strengths and weaknesses of these two strategies, the ideal approach should combine both: decomposing problems into several coarse-grained subproblems (HTS) and applying the VTS to each subproblem. Thus, researching how to improve vertical thinking capability and understanding why horizontal thinking strategy can significantly enhance model reasoning ability are both crucial.
@@
In this work, we investigate the performance of Transformer models on a symbolic multi-step reasoning dataset. Our work aims to uncover these mechanisms and provide insights into how Transformer processes and integrates logical information across multiple layers to perform multi-step reasoning, which can help develop more effective strategies for improving their multi-step reasoning abilities. Specifically, we found that Transformers utilize a Buffer Mechanism when engaging in complex reasoning. The model stores different intermediate results in separate buffers, allowing for quick retrieval as needed. We elaborate on how the model leverages this buffer mechanism for vertical thinking and we explain why horizontal thinking strategy can significantly enhance the model’s multi-step reasoning capability from the perspective of the buffer mechanism. Finally, based on this understanding, we propose a method to enhance the model’s reasoning abilities, resulting in a 75% reduction in the cost of generalization for the GPT-2 model on the question-answering dataset PrOntoQA (Saparov & He, 2022). The concept of “buffer" or similar concepts has also been mentioned in other works (Reddy, 2023; Bietti et al., 2024; Elhage et al., 2021). Our work provides a detailed description of the concept of buffer and applies this mechanism to enhance model performance.
@@
  • We discover the buffer mechanism employed by language models during the reasoning process and provided a detailed analysis of the model’s internal thinking process for vertical and horizontal thinking from the perspective of the buffer.
  • We propose a method to enhance the model’s reasoning capability, significantly improving data utilization efficiency in logical reasoning datasets.
Our research deepens the understanding of the reasoning mechanisms in Transformer models and provides new perspectives for further enhancing their reasoning capability. The insights gained from this study can contribute to the design of more efficient reasoning models and the exploration of reasoning mechanisms in general artificial intelligence systems.
@@ 2 Reasoning Dataset and Transformer Model
Dataset. To understand the mechanism of multi-step reasoning in Transformers, we design an abstract symbolic multi-step reasoning task. As shown in Fig. 2, reasoning chains are serialized into a sequence. Every two tokens in the sentence represent a reasoning relation. The last token is the reasoning start token, and the label is the reasoning result with a fixed-step reasoning starting from the starting point.
Figure 2: Illustration of the dataset and Transformer structure.
Transformer Model. We employ a decoder-only Transformer. Given an input sequence ${\bm{X}}^{\mathrm{in}}\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the dictionary size, the model first applies an embedding layer (target embedding and position embedding) to obtain the input representation ${\bm{X}}^{(1)}={\bm{X}}_{tgt}+{\bm{X}}_{pos}\in\mathbb{R}^{n\times d_{m}}$. The single-head attention in each layer is computed as follows:
$\mathcal{A}^{(l)}({\bm{X}})=\sigma\left(\frac{\mathrm{mask}({\bm{X}}{\bm{W}}^{q(l)}{\bm{W}}^{k(l),\mathsf{T}}{\bm{X}}^{\mathsf{T}})}{\sqrt{d_{k}}}\right),\quad{\bm{X}}^{\mathrm{qkv}(l)}=\mathcal{A}^{(l)}(\bar{{\bm{X}}}^{(l)})\bar{{\bm{X}}}^{(l)}{\bm{W}}^{v(l)}{\bm{W}}^{o(l)},$
@@
${\bm{X}}^{\mathrm{ao}(l)}={\bm{X}}^{(l)}+{\bm{X}}^{\mathrm{qkv}(l)},\quad{\bm{X}}^{(l+1)}=f^{(l)}(\bar{{\bm{X}}}^{\mathrm{ao}(l)})+{\bm{X}}^{\mathrm{ao}(l)},$
@@
${\bm{Y}}=argmax(\sigma(\bar{{\bm{X}}}^{(L)}{\bm{W}}^{p}))\in\mathbb{R}^{n}.$ (1)
@@
$\texttt{x}_{\texttt{2i}}-\texttt{x}_{\texttt{2i-1}}\ \ (\text{mod }m)\in G.$ (2)
@@
$\texttt{x}_{\texttt{2i}}-\texttt{x}_{\texttt{2i-1}}\ \ (\text{mod }m)\in\{1,\cdots,m\}\backslash G,$ (3)
@@
3 Buffer Mechanism and Vertical Thinking Strategy
In this section, we examine the mechanism by which the Transformer model employs vertical thinking strategy for multi-step reasoning. We discovered that when the number of layers exceeds the number of reasoning steps, the Transformer’s logic circuit of vertical thinking can be depicted in Fig. 3. The first layer facilitates the fusion of token information at odd and even positions. In each subsequent layer, which we refer to as the reasoning layer, the token at the last position attends to the token carrying the next reasoning result based on existing reasoning result, thereby accomplishing one-step reasoning.
Figure 3: Illustration of how the Transformer model employs vertical thinking strategy for multi-step reasoning task.
@@ 3 Buffer Mechanism and Vertical Thinking Strategy
We observe that each token in Fig. 3 stores various information in various forms (projected by different matrices ${\bm{W}}^{vo(l)}$), which leads to a natural question: how does the attention layer effectively retrieve useful information while avoiding interference from others? To address this question, we introduce the following lemma (the proof can be found in Appendix B):
Lemma 1.
Suppose token ${\bm{x}}=\sum_{i=1}^{n}{\bm{a}}_{i}{\bm{W}}_{i}\in\mathbb{R}^{d_{m}}$ and token ${\bm{y}}=\sum_{i=1}^{n}{\bm{b}}_{i}{\bm{W}}_{i}\in\mathbb{R}^{d_{m}}$, where ${\bm{a}}_{i},{\bm{b}}_{i}\in\mathbb{R}^{d_{m}}$, ${\bm{W}}_{i}\in\mathbb{R}^{d_{m}\times d_{m}},\ i=1,2,\cdots,n$. Each element of $\{{\bm{a}}_{i}\}_{i=1}^{n}$, $\{{\bm{b}}_{i}\}_{i=1}^{n}$ and $\{{\bm{W}}_{i}\}_{i=1}^{n}$ follows a normal distribution $\mathcal{N}(0,1/d_{m})$ and independent to others. Then, we have:
@@
$h^{(l)}({\bm{X}})$ $=({\bm{X}}{\bm{W}}^{vo(l-1)}){\bm{W}}^{q(l)}{\bm{W}}^{k(l),\mathsf{T}}({\bm{X}}{\bm{W}}^{vo(0)})^{\mathsf{T}}$ (6)
$={\bm{X}}{\bm{W}}^{vo(l-1)}{\bm{W}}^{qk(l)}{\bm{W}}^{vo(0),\mathsf{T}}{\bm{X}}^{\mathsf{T}}\triangleq{\bm{X}}\,\text{Ker}^{(l)}\,{\bm{X}}^{\mathsf{T}},\ l\geq 2,$ (7)
$h^{(1)}({\bm{X}})$ $={\bm{X}}{\bm{W}}^{qk(1)}{\bm{W}}^{vo(0),\mathsf{T}}{\bm{X}}^{\mathsf{T}}\triangleq{\bm{X}}\,\text{Ker}^{(1)}\,{\bm{X}}^{\mathsf{T}}.$ (8)
@@
where we temporarily ignore the effects introduced by the feedforward layers only in our analysis; a detailed version that includes the feedforward layers can be found in Appendix C. When same-token matching happens, the following should hold:
$\text{Ker}^{(l)}$ $={\bm{W}}^{vo(l-1)}{\bm{W}}^{qk(l)}{\bm{W}}^{vo(0),\mathsf{T}}\approx{\bm{I}},\ l\geq 2$ (9)
$\text{Ker}^{(1)}$ $={\bm{W}}^{qk(1)}{\bm{W}}^{vo(0),\mathsf{T}}\approx{\bm{I}},$ (10)
$h^{(l)}({\bm{X}}_{tgt})$ $\approx{\bm{X}}_{tgt}{\bm{X}}_{tgt}^{T}={\bm{I}}+\mathcal{O}\left(\frac{1}{\sqrt{d_{m}}}\right),$ (11)
where we assume ${\bm{X}}_{tgt}$ remaining a random matrix with minimal changes during training. Eq.(9), Eq.(10), and Eq.(11) can explain why the model exhibits out-of-distribution generalization capability. Since the embedding vectors of $\texttt{token}_{\texttt{OOD}}$ are untrained random vectors when the model’s weights satisfy Eq.(9) and Eq.(10), then Eq.(11) automatically holds. This indicates that the model has learned the underlying patterns in the data.
3.1 Experiments
We utilized a 3-layer, single-head Transformer model to learn from a two-step reasoning dataset. Specific Experimental settings can be found in Appendix A. After training, the Transformer exhibits generalization capability in in-distribution (100% accuracy) and out-of-distribution data (over 82% accuracy).
Fig. 4 presents the experimental validation of Eq.(9) $\sim$ Eq.(11). We visualize the computed values of $h^{(1)}({\bm{X}}_{tgt})$, $h^{(2)}({\bm{X}}_{tgt})$, $\text{Ker}^{(1)}$, and $\text{Ker}^{(2)}$ with the model parameters after training. Notably, in the $\texttt{token}_{\texttt{OOD}}$ region, $\texttt{token}_{\texttt{OOD}}$ and $h^{(2)}({\bm{X}}_{tgt})$ exhibit a diagonal structure similar to an identity matrix ${\bm{I}}$, ensuring the model’s OOD generalization capability.
Figure 4: (a) (b) Heatmap of $h^{(1)}({\bm{X}}_{tgt})$ and $h^{(2)}({\bm{X}}_{tgt})$. (c) (d) Heatmap of $\text{Ker}^{(1)}$ and $\text{Ker}^{(2)}$. For better presentation, we only plot $\text{Ker}^{(1,2)}_{[:60,:60]}$ (the actual dimension of $\text{Ker}^{(1,2)}$ is $400\times 400$).
To establish a strong correlation between the model’s OOD generalization and its ability to use the buffer mechanism for same-token matching, we define a metric for this capability, the matching score:
Matching $\text{Matching Score}(h^{{(l)}})=\frac{1}{n}\mathbb{E}_{{\bm{X}}}\text{Trace}(\sigma(h^{(l)}({\bm{X}}))),$ (12)
where ${\bm{X}}\in{\mathbb{R}}^{n\times d_{m}}$ is sampled from $\texttt{token}_{\texttt{ID}}$ or $\texttt{token}_{\texttt{OOD}}$ for expectation $\mathbb{E}_{{\bm{X}}}$, and $\sigma$ takes the softmax operation. The level of diagonalization of $\text{Ker}^{(l)}$ serves as the intrinsic driver for achieving same-token matching; thus, we also define a metric for this diagonalization level, the kernel score:
Kernel Score($\text{Kernel Score}(\text{Ker}^{(l)})=\frac{n\cdot\text{Trace}(\text{Ker}^{(l)})}{\sum_{i,j}\text{Ker}^{(l)}_{i,j}}.$ (13)
Fig. 5 illustrates the accuracy curve and the dynamic changes in the matching score and kernel score during training. It is observed that the increase in the model’s ID and OOD generalization coincide with the increases in the model’s matching score and kernel score. Based on the above experimental analysis, we conclude that the Transformer model indeed employs the buffer mechanism for vertical thinking.
Figure 5: (a) Accuracy curve during training. (b) The dynamic evolution of the model’s matching score. The red and blue lines represent the matching scores for the first and second layers, respectively. Solid and dashed lines indicate the matching scores when ${\bm{X}}$ is drawn from $\texttt{token}_{\texttt{ID}}$ and $\texttt{token}_{\texttt{OOD}}$, respectively. (c) The kernel scores for the first (red) and second (blue) layers.
According to the Buffer mechanism, in theory, by setting the model’s weights in the following way, an $L$-layer model can achieve $(L-1)$-step reasoning without the extra training process:
where $\{{\bm{W}}^{vo(l)}\}_{l=1}^{L}$ are set as random matrices and the projection layer satisfies ${\bm{W}}^{p}={\bm{W}}^{vo(L),\mathsf{T}}{\bm{W}}^{\text{emb},\mathsf{T}}$. Under this configuration, an 8-layer Transformer model can achieve 7-step reasoning (See Appendix C).
Another natural question is whether the model can still employ this mode of reasoning when the number of layers $L$ is less than the number of reasoning steps. In fact, in this scenario, the model exhibits the ability to engage in parallel thinking. Specifically, the model can partition $2^{L}$ buffers through $\{\prod_{l\in J}{\bm{W}}^{vo(l)}|J\subset\{0,1,\cdots,L-1\}\}$ since each $\prod_{l\in J}{\bm{W}}^{vo(l)}$ can be regarded as a new random matrix with the same distribution as ${\bm{W}}^{vo}$. We discuss this issue in detail in Appendix D.
4 Understanding Horizontal Thinking Strategy With Buffer Mechanism

### v3

# Understanding the Language Model to Solve the Symbolic Multi-Step Reasoning Problem from the Perspective of Buffer Mechanism

@@ Abstract
Large language models have consistently struggled with complex reasoning tasks, such as mathematical problem-solving. Investigating the internal reasoning mechanisms of these models can help us design better model architectures and training strategies, ultimately enhancing their reasoning capability. In this study, we constructed a symbolic multi-step reasoning task to investigate the information propagation mechanisms in Transformer models when solving the task through direct answering and Chain-of-Thought (CoT) reasoning. We introduced the concept of buffer mechanism: the model stores various information in distinct buffers and selectively extracts it through the query-key matrix. We proposed a random matrix-based algorithm to enhance the model’s reasoning ability. This algorithm introduces only 132 trainable parameters, yet leads to significant performance improvements on 7 multi-step reasoning datasets, including PrOntoQA, LogicAsker, and LogicInference. These findings provide new insights into understanding the large language models.
Understanding the Language Model to Solve the Symbolic Multi-Step Reasoning Problem from the Perspective of Buffer Mechanism
Zhiwei Wang1,2, Yunji Wang1,2, Zhongwang Zhang1,2, Zhangchen Zhou1,2, Hui Jin4, Tianyang Hu3, Jiacheng Sun4, Zhenguo Li 4, Yaoyu Zhang1,2, Zhi-Qin John Xu1,2,5 1Institute of Natural Sciences, MOE-LSC, Shanghai Jiao Tong University 2School of Mathematical Sciences, Shanghai Jiao Tong University 3The Chinese University of Hong Kong, Shenzhen 4Huawei Noah’s Ark Lab 5Key Laboratory of Marine Intelligent Equipment and System, Ministry of Education Correspondence: xuzhiqin@sjtu.edu.cn
@@ 1 Introduction
In recent years, LLMs have emerged and demonstrated remarkable capability across various tasks (Vaswani et al., 2017; Liu et al., 2018; Devlin et al., 2018; Radford et al., 2019; Touvron et al., 2023; OpenAI, 2023; Liu et al., 2024; Guo et al., 2024, 2025). These models have shown impressive in-context learning abilities (Brown et al., 2020; Dong et al., 2022; Garg et al., 2022) and have been applied to logical reasoning problems, such as matching top human contestants at the International Mathematical Olympiad (IMO) level (Trinh et al., 2024) and solving math problems (Davies et al., 2021). However, even the most advanced models still struggle with complex reasoning tasks. To truly enhance the reasoning capability of LLMs, it is crucial to investigate their intrinsic mechanisms.
[…]
Figure 1: The interaction results of multi-step reasoning tasks with large models. We tested each model 25 times, and when the number of reasoning steps exceeded three, these models exhibited random guessing. However, when we allowed the models to use CoT prompting (by removing “directly" and “only return the answer"), all models achieved 100% accuracy. Detailed interaction results are provided in the Appendix I.
@@
Large models employ two primary strategies for logical reasoning. The first, known as the Vertical Thinking Strategy (VTS), outputs reasoning results in a single forward based on their inherent structure. This approach is efficient but demands a high level of intelligence. As shown in Fig. 1, current large models exhibit significant limitations in their vertical thinking capability. Another relatively less efficient approach is the (HTS), such as Chain of Thought (CoT) (Wei et al., 2022; Kojima et al., 2022) and Tree of Thought (ToT) (Yao et al., 2024), Diagram of Thought (DoT) (Zhang et al., 2024a). This strategy substantially enhances the model’s reasoning performance. All models can produce the correct answers for tasks depicted in Fig. 1 with the help of CoT. Considering the strengths and weaknesses of these two strategies, the ideal approach should combine both: decomposing problems into several coarse-grained subproblems (HTS) and applying the VTS to each subproblem. Thus, researching how to improve vertical thinking capability and understanding why the horizontal thinking strategy can significantly enhance model reasoning ability are both crucial.
@@
In this work, we investigate the performance of Transformer models on a symbolic multi-step reasoning dataset. Our work aims to uncover these mechanisms and provide insights into how Transformers process and integrate logical information across multiple layers to perform multi-step reasoning, which can help develop more effective strategies for improving their multi-step reasoning abilities. Specifically, we found that Transformers utilize a Buffer Mechanism when engaging in symbolic multi-step reasoning tasks. The model stores different intermediate results in separate buffers, allowing for quick retrieval as needed. We elaborate on how the model leverages this buffer mechanism for vertical thinking, and we explain why the horizontal thinking strategy can significantly enhance the model’s multi-step reasoning capability from the perspective of the buffer mechanism. Finally, based on this understanding, we propose a method to improve the model’s reasoning abilities, leading to significant performance improvements for the GPT-2 model on 7 multi-step reasoning datasets, including Clutrr (Sinha et al., 2019), RuleTaker (Clark et al., 2020), StepGame (Shi et al., 2022), LogicInference (Ontanon et al., 2022), LogicAsker (Wan et al., 2024), PararulePlus (Bao et al., 2022), and PrOntoQA (Saparov and He, 2022).
The concept of “buffer" or similar concepts has also been mentioned in other works (Reddy, 2023; Bietti et al., 2024; Elhage et al., 2021). Our work provides a detailed description of the concept of buffer and applies this mechanism to enhance model performance.
@@
  • We propose a buffer mechanism and found evidence that supports such a mechanism being employed by language models during the reasoning process in symbolic multi-step reasoning tasks and provide a detailed analysis of the model’s internal thinking process for vertical and horizontal thinking from the perspective of the buffer.
  • We propose a method to enhance the model’s reasoning capability, improving data utilization efficiency in 7 logical reasoning datasets.
Our research deepens the understanding of the reasoning mechanisms in Transformer models.
@@ 2 Reasoning Dataset and Transformer Model
Dataset. To understand the mechanism of multi-step reasoning in Transformers, we design an abstract symbolic multi-step reasoning task. As shown in Fig. 2, reasoning chains are serialized into a sequence. Every two tokens in the sentence represent a reasoning relation. The last token is the reasoning start token, and the label is the result with a fixed-step reasoning starting from the starting point.
Figure 2: Illustration of the dataset. We investigate reasoning chains composed of digital tokens. In a serialized representation, each pair of adjacent tokens (represented by the same color) forms a reasoning relation within the reasoning chain. The order of the reasoning relations is random, thus a single reasoning chain can correspond to various serialized representations. The model input consists of the serialized representation along with a reasoning start token. The label is the result of performing a fixed number of reasoning steps starting from this start token. This figure illustrates a 2-step reasoning task.
Transformer Model. We employ a decoder-only Transformer. Given an input sequence ${\bm{X}}^{\mathrm{in}}\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the dictionary size, the model first applies an embedding layer to obtain the input representation ${\bm{X}}^{(1)}={\bm{X}}_{tgt}+{\bm{X}}_{pos}\in\mathbb{R}^{n\times d_{m}}$, where ${\bm{X}}_{tgt}$ and ${\bm{X}}_{pos}$ denote the token and positional embeddings, respectively, and $d_{m}$ is the hidden dimension. The single-head attention in each layer is computed as follows:
$\mathcal{A}^{(l)}({\bm{X}})$ $=\sigma\left(\frac{\mathrm{mask}({\bm{X}}{\bm{W}}^{q(l)}{\bm{W}}^{k(l),\mathsf{T}}{\bm{X}}^{\mathsf{T}})}{\sqrt{d_{k}}}\right),$
$\quad{\bm{X}}^{\mathrm{qkv}(l)}$ $=\mathcal{A}^{(l)}(\bar{{\bm{X}}}^{(l)})\bar{{\bm{X}}}^{(l)}{\bm{W}}^{v(l)}{\bm{W}}^{o(l)},$
@@
${\bm{X}}^{\mathrm{ao}(l)}$ $={\bm{X}}^{(l)}+{\bm{X}}^{\mathrm{qkv}(l)},$
${\bm{X}}^{(l+1)}$ $=f^{(l)}(\bar{{\bm{X}}}^{\mathrm{ao}(l)})+{\bm{X}}^{\mathrm{ao}(l)},$
@@
${\bm{Y}}=argmax(\sigma(\bar{{\bm{X}}}^{(L)}{\bm{W}}^{p}))\in\mathbb{R}^{n}.$
@@
$\texttt{x}_{\texttt{2i}}-\texttt{x}_{\texttt{2i-1}}\ \ (\text{mod }m)\in G.$
@@
$\G,\texttt{x}_{\texttt{2i}}-\texttt{x}_{\texttt{2i-1}}\ \ (\text{mod }m)\in\{1,\cdots,m\}\backslash$ G,
@@
Figure 3: Illustration of how the Transformer employs the vertical and horizontal thinking strategies to solve the multi-step reasoning task. The information crucial for the layer’s information transfer is highlighted. The first attention pairs tokens at odd positions with those at even positions. In the Vertical Thinking Strategy (VTS), each layer conveys its newly derived reasoning result to the final token. In the Horizontal Thinking Strategy (HTS), all information transfer is accomplished within a single layer.
3 Vertical and Horizontal Thinking Strategy
We illustrate how the Transformer model employs vertical or horizontal thinking strategies for multi-step reasoning. Fig. 3 depicts the schematic mechanisms by which these two strategies execute multi-step inference. The critical logic circuits highlighted in this figure were identified through causal intervention experiments (Feng and Steinhardt, 2023; Meng et al., 2022; Vig et al., 2020; Wang et al., 2024a) (Appendix H). From the figure, one can discern the core distinction between VTS and HTS. Under VTS, in every layer beyond the first—termed a “reasoning layer”—the final token attends to the token carrying the next inference result, conditional on the existing result, thereby effecting a single reasoning step. In contrast, HTS accomplishes all logical inference within a single layer.
This observation naturally prompts a key question: irrespective of whether VTS or HTS is employed, each node in the Transformer stores multiple pieces of information. For instance, the red node in Layer 1 encodes not only [a] but also [b]. How, then, does the model consolidate diverse information within a single node and accurately extract and utilize the relevant subset? Addressing this question lies at the heart of our investigation into multi-step reasoning.
4 Buffer Mechanism
To address the above problem, we propose a mechanism employed by the Transformer, which we refer to as the “buffer mechanism”.
@@ 4 Buffer Mechanism
Below, we introduce the buffer mechanism within the VTS framework. All theoretical results are supported by experimental validation (Section 5). We observe that each token in Fig. 3(a) stores various information in various forms (projected by different matrices ${\bm{W}}^{vo(l)}$), which leads to a natural question: how does the attention layer effectively retrieve useful information while avoiding interference from others? To address this question, we introduce the following lemma (the proof can be found in Appendix C):
Lemma 4.1.
Suppose token ${\bm{x}}=\sum_{i=1}^{n}{\bm{a}}_{i}{\bm{W}}_{i}\in\mathbb{R}^{d_{m}}$ and token ${\bm{y}}=\sum_{i=1}^{n}{\bm{b}}_{i}{\bm{W}}_{i}\in\mathbb{R}^{d_{m}}$, where ${\bm{a}}_{i},{\bm{b}}_{i}\in\mathbb{R}^{d_{m}}$, ${\bm{W}}_{i}\in\mathbb{R}^{d_{m}\times d_{m}},\ i=1,2,\cdots,n$. Each element of $\{{\bm{a}}_{i}\}_{i=1}^{n}$, $\{{\bm{b}}_{i}\}_{i=1}^{n}$ and $\{{\bm{W}}_{i}\}_{i=1}^{n}$ follows $\mathcal{N}(0,1/d_{m})$ and independent to others. Then, we have:
${\bm{x}}{\bm{W}}_{i}^{\mathsf{T}}$ $={\bm{a}}_{i}+\mathcal{O}\left(\sqrt{\frac{n}{d_{m}}}\right),$
${\bm{y}}{\bm{W}}_{j}^{\mathsf{T}}$ $={\bm{b}}_{j}+\mathcal{O}\left(\sqrt{\frac{n}{d_{m}}}\right),$
${\bm{x}}{\bm{W}}_{i}^{\mathsf{T}}{\bm{W}}_{j}{\bm{y}}^{\mathsf{T}}={\bm{a}}_{i}{\bm{b}}_{j}^{\mathsf{T}}+\mathcal{O}\left(\frac{n}{\sqrt{d_{m}}}\right).$
@@
$h^{(1)}$ $({\bm{X}})={\bm{X}}{\bm{W}}^{qk(1)}{\bm{W}}^{vo(0),\mathsf{T}}{\bm{X}}^{\mathsf{T}}\triangleq{\bm{X}}\,\text{Ker}^{(1)}\,{\bm{X}}^{\mathsf{T}},$
$h^{(l)}$ $({\bm{X}})$ (1)
$=({\bm{X}}{\bm{W}}^{vo(l-1)}){\bm{W}}^{q(l)}{\bm{W}}^{k(l),\mathsf{T}}({\bm{X}}{\bm{W}}^{vo(0)})^{\mathsf{T}}$
$={\bm{X}}{\bm{W}}^{vo(l-1)}{\bm{W}}^{qk(l)}{\bm{W}}^{vo(0),\mathsf{T}}{\bm{X}}^{\mathsf{T}}$
$\triangleq{\bm{X}}\,\text{Ker}^{(l)}\,{\bm{X}}^{\mathsf{T}},\ l\geq 2.$ (2)
where
$\text{Ker}^{(1)}$ $={\bm{W}}^{qk(1)}{\bm{W}}^{vo(0),\mathsf{T}},$
$\text{Ker}^{(l)}$ $={\bm{W}}^{vo(l-1)}{\bm{W}}^{qk(l)}{\bm{W}}^{vo(0),\mathsf{T}},\ l\geq 2.$ (3)
@@
We temporarily ignore the effects introduced by the feedforward layers only in our analysis; a detailed version that includes the feedforward layers can be found in Appendix D.
To achieve the same-token matching, it is sufficient for $\text{Ker}^{(l)}\approx I$, in which case
$h^{(l)}({\bm{X}}_{tgt})\approx{\bm{X}}_{tgt}{\bm{X}}_{tgt}^{\mathsf{T}}={\bm{I}}+\mathcal{O}\left(\frac{1}{\sqrt{d_{m}}}\right).$ (4)
Eq.(4) means that the attention of all tokens is focused on themselves.
Furthermore, a much more remarkable observation is that the same-token matching is independent of the specific value of ${\bm{X}}_{tgt}$. For example, for ${\bm{X}}_{tgt,\text{OOD}}$ sampled from the untrained random vectors $\texttt{token}_{\texttt{OOD}}$, $h^{(l)}({\bm{X}}_{tgt,\text{OOD}})\approx{\bm{I}}$ still holds. Therefore, when $\text{Ker}^{(l)}\approx I$ holds, the model exhibits OOD generalization capability.